OpenStudy (anonymous):

sorry. but I need help. Elena's daughter sold 102 boxes of Girl Scout cookies. She was told that this amount represented 5% of the troop's sales. How many boxes did the troop sell?

OpenStudy (anonymous):

5% of the troop's sales is equal to 102, you need to find 100%.

OpenStudy (anonymous):

Try to think about it. It's not difficult at all.

OpenStudy (anonymous):

so do i make a proportion?

OpenStudy (dumbcow):

you could 5/100 = 102/x

OpenStudy (dumbcow):

or think about it...how many times does 5 go into 100?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Exactly!! then how many boxes did the troop sell?

OpenStudy (anonymous):

2040 ?

OpenStudy (anonymous):

102=0.05boxes boxes = 102/0.05 ..........=2040boxes